find the slope and y intercept
is the intercept wher ethe line crosses the asis
3x+5y+10=0

Answer by Earlsdon(6103)

One way you can do this is to get the equation into the "slope-intercept" form: y = mx + b where m is the slope and b is the y-intercept.
3x+5y+10 = 0 Subtract 3x from both sides of the equation.
5y+10 = -3x Now subtract 10 from both sides.
5y = -3x-10 Finally, divide both sides by 5.
y = (-3/5)x-2 Compare this with the slope-intercept form.
y = mx+b and you can see that m = -3/5 This is the slope, and b = -2 This is the y-intercept which is where the line crosses the y-axis.